Grade 11General Physics

Rain falls with a speed of 10m/s, at an angle of 53 degrees with the vertical. If a man observes rain falling at an angle of 37 degrees, what is his speed?

Saurabh
8 Years agoGrade 11

1 Answer

DIWAKAR PANDEY
8 Years ago
Rain fall with velocity of 10 m/s and 53° with vertical. Man observe it 37° so by resolution it is 10 sin53= v (v=speed of man)By using it V=10×4/5 V=8 m/s